OpenStudy (anonymous):

Find the equation of the line that is parallel to the line formed by the equation 3x + 2y = -20 and passes through the point (-24, 12).

OpenStudy (ddcamp):

It will be easier to find the slope if you put the equation in slope intercept (y=mx+b) form by solving for y. Once you have the slope, you can plug it into point slope (y-y₁=m(x-x₁)) form by plugging in (-24,12) for (x₁, y₁)

OpenStudy (anonymous):

Since line is || to 3x + 2y = -20 i.e. 2y=-3x-20 i.e. y= -3/2x-10 comparing it with y=mx+c we find slope m= -3/2 Since line passes through the point (-24, 12) therefore \[x_{1} =-24 \] \[ y_{1} =12\] The eq of line in slope point form is given as \[y-y_{1} =m(x-x_{1})\] plugging in the related values in the above eq we find \[y-12= \frac{-3}{2}(x+24)\] i.e. \[y-12= \frac{-3}{2}x-36 \rightarrow y= \frac{-3}{2}x-36 + 12\] i.e. \[y= \frac{-3}{2}x-24 \rightarrow 2y = -3x-48 \rightarrow 3x+2y =-48\] Thus 3x+2y =-48 is the required eq of the given line